OpenStudy (anonymous):

My question is to find the median of the following data set: Books read: 0 - 2 3 - 5 6 - 8 9 - 11 12 - 14 Number of people: 20 60 50 120 15

OpenStudy (anonymous):

I know what median is and how to find it... but the data set is confusing to me since it isn't exact :/

OpenStudy (jdoe0001):

the data doesn't have to be "exact", it just have to be the same for each set to make the set and correlate

OpenStudy (jdoe0001):

sort them 1st, then take it from there :)

OpenStudy (anonymous):

Books read will be 6-8 and number of people will be 50 For me :/

OpenStudy (anonymous):

Yes, I know. But how am I supposed to sort numbers that aren't even there? Like "0-2" How do I sort THAT out

OpenStudy (anonymous):

@xoya what....? Why did you restate what I posted?

OpenStudy (jdoe0001):

well, 0-2 is just 1 item of a set

OpenStudy (anonymous):

I gave 0-2 as an example. you have 0 - 2 3 - 5 6 - 8 9 - 11 12 - 14. HOW do I lay those out? I know 20 people read 0-2 books but I don't know if they read ONE book, TWO books, or ZERO books

OpenStudy (anonymous):

LOL leavee :(

OpenStudy (jdoe0001):

the data set can contain individuals, or groups in this case your data set 1 is a group while the data set 2 is just an individual still each of the items are treated as singular in themselves, inside the data set

OpenStudy (anonymous):

@ivettef365 @MathLegend @Mertsj @mathstudent55

jimthompson5910 (jim_thompson5910):

how many people are there total

OpenStudy (anonymous):

265 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so the median will be the person who is at position 133 since 265/2 = 132 remainder 1 because you have an odd number of people, the median will be a number in this list 132 people will be below the median 132 people will be above the median the median will be between the first 132 people and the last 132 people

jimthompson5910 (jim_thompson5910):

notice how 20+60 = 80, which is less than 132 and how 20+60+50 = 130, still too low

jimthompson5910 (jim_thompson5910):

you just need 3 more people, and that lands you in the 9 - 11 range

OpenStudy (anonymous):

That honestly makes no sense... I don't know how to lay my information out, that's all i need... I know how to find the median but my data set is confusing. do i lay it out as: 20, 20, 60, 60, 60, 60, 60, 50, 50, 50, 50, 50, 50, 50, 50, 120, 120, 120..... ETC

jimthompson5910 (jim_thompson5910):

ok, well when they say there are 20 people in the 0-2 range, it means there are 20 people who read 0 books, 1 book or 2 books

OpenStudy (anonymous):

Mhm

jimthompson5910 (jim_thompson5910):

so the first row of 0-2 20 *could* translate to this list 0,0,0,0,0,0,1,1,1,1,1,1,2,2,2,2,2,2,2,2,2

jimthompson5910 (jim_thompson5910):

or it could be this 1,1,1,1,1,1,1,1,1,1,2,2,2,2,2,2,2,2,2,2

jimthompson5910 (jim_thompson5910):

all we know is that there are 20 people that read between 0 and 2 books

OpenStudy (anonymous):

I think i'm better off if I call my teacher....

jimthompson5910 (jim_thompson5910):

The first 20 read between 0 and 2 books the next 60 read between 3 and 5 books so 20+60 = 80 read between 0 and 5 books

jimthompson5910 (jim_thompson5910):

the next 50 read between 6 and 8 books so 20+60+50 = 130 people read between 0 and 8 books

OpenStudy (anonymous):

Thanks, but I'm just going to call my teacher about it. I only partially understand what you're telling me

jimthompson5910 (jim_thompson5910):

if you have them sorted by the number of books read, then the first 132 people will be below the median so far, we're at 130 people

jimthompson5910 (jim_thompson5910):

ok fine

OpenStudy (anonymous):

Thanks for trying~!